OpenStudy (anonymous):

Test the series for convergence or divergence. \[\sum_{2}^{\infty}\left( \left( -1 \right)^{n}\ln n \right)\div \left( n \right)\]

OpenStudy (anonymous):

. \[\sum_{2}^{\infty}\left( \left( -1 \right)^{n}\ln n \right)\div \left( n \right)\]